OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 14x + 18y = -54 -4x – 14y = 42

OpenStudy (saifoo.khan):

JB, give it a try man!

OpenStudy (anonymous):

I'm justin bieber

OpenStudy (anonymous):

\[\{x=0,y=-3\} \]

OpenStudy (saifoo.khan):

Justin Bieber the great: http://www.youtube.com/watch?v=zqZ42LmDou4

OpenStudy (anonymous):

14x + 18y = -54 ----(I) -4x – 14y = 42 ----(II) multiply eqn(I) with 2 and eqn(II) with 7, we get 28x + 36y = -108 -28x – 98y = 294 add the above 2 eqns we get -62y=186 so y=-3 put y=-3 in any one of the given eqns -4x – 14y = 42 -4x – 14*(-3) = 42 -4x+42=42 -4x=0 x=0 so y=-3 and x=0